# One Model to Rule them All: Towards Universal Segmentation for Medical Images with Text Prompts

### v2

@@ 2.2 Ablation Study
As will be illustrated in Sec. 6, to build an universal segmentation model driven by text prompts, we devise the visual encoder and decoder to process diverse medical images and perform segmentation; Meanwhile, domain knowledge is injected into the text encoder to provide precise prompts for the segmentation target, i.e., the encoding of terminology. In this section, we conduct experiments and discuss the effect of different visual backbones and domain knowledge in SAT. To save computational cost, all the experiment in this section are conduct on a subset of SAT-DS, termed ast SAT-DS-Nano. It includes 49 datasets, 13,303 images, 151,461 annotations and 429 classes.
@@ 2.2.1 Effect of Visual Backbone
Extensive studies have been done discussing the applicability of different visual backbones across segmentation tasks. However, it remains unexplored on which visual backbone suits the need of universal segmentation. In addition to the ConvNet-based U-Net, we consider two alternative backbones, namely, SwinUNETR [31] and U-Mamba [29] to medical segmentation. We configure three SAT-Nano variants with SwinUNETR(107M), U-Mamba(114M) and U-Net(110M) in comparable size: For SwinUNETR, we use the same hypeparameters as in official implementation [30]; For U-Mamba, we refer to the official implementation [67], reserve configuration of the U-Net in SAT-Nano, and only add mamba layers at the end of last 3 blocks of the U-Net encoder.
To avoid repeating multimodal knowledge injection on each visual backbones, we use MedCPT [42] for all these varients as text encoder, without loss of generality. Like the text encoder we pre-trained in Sec. 6.2, MedCPT is a text encoder trained on 255 million in-house user click logs from PubMed [91], and shows state-of-the-art performance on various biomedical language tasks. We expect it to possess domain knowledge and embed terminologies properly as text prompts. Thus, we denote these variants as U-NET-CPT, SwinUNETR-CPT and U-Mamba-CPT. We demonstrate the region-wise evaluation results in Fig. 6. U-Net-CPT outperforms U-Mamba-CPT slightly on both DSC (0.35) and NSD (0.22) scores averaged over all classes. However, both U-Net-CPT and U-Mamba-CPT exceed SwinUNETR-CPT by a significant margin. These observations confirms our choice of U-Net as the visual backbone of SAT. For more detailed comparison on each dataset and class, we refer the reader to Table LABEL:tab:visual_ab_dataset_results, Table 9 and Table 10.
@@ 2.2.2 Effect of Text Encoder
For an universal segmentation model with text as prompts, the neural representations of anatomical structures and lesions serve as the guidance for the segmentation model. And thus, the choice of text encoder should be carefully studied. In this section, we investigate the impact of domain knowledge on building a text encoder for medical universal segmentation task. Specifically, we train three SAT-Nano models with three representative text encoders: the text encoder pre-trained on our multimodal medical knowledge graph, as illustrated in Sec. 6.2; the state-of-the-art text encoder for various medical language tasks, i.e., MedCPT; And BERT-Base, a prevalent text encoder in natural language processing but not specifically finetuned on medical corpus. For all the variants, we use U-Net as visual backbone, and denote them as U-Net-Ours, U-Net-CPT, U-Net-BB.
@@
Region-wise evaluation. As shown in Fig. 7, the performance of U-Net-CPT and U-Net-BB are close, with U-Net-CPT slightly outperfoms U-Net-BB on both DSC (+0.43) and NSD (+0.52) scores averaged over all classes, suggesting that finetuned on medical corpus, MedCPT might hold minor advantage. By contrast, U-Net-Ours surpasses U-Net-CPT consistently on all regions and lesions, leading notable margins on both DSC (+1.54) and NSD (+2.65) scores after average over all classes. This demonstrates the effectiveness of multimodal knowledge injection for training SAT. For more detailed comparison on each dataset and class, we refer the reader to Table LABEL:tab:text_ab_dataset_results, Table 6 and Table 7.
@@
Concept-to-definition retrieval. To demonstrates advantage our proposed knowledge injection on capturing the critical features of medical terminologies, we evaluate the three text encoders on concept to definition retrieval task, with the human anatomy knowledge we collect, totaling 6,502 anatomy concepts and definition pairs. We find that the recall at 1 (R1) for BERT-Base is merely 0.08%, suggesting it can hardly understand these anatomy concepts or possess almost no domain knowledge; The R1 for MedCPT is 11.19%. Though this is a significant improvement over BERT-Base, MedCPT still struggles to distinguish these concepts; By contrast, our proposed text encoder get 99.18% R1, indicating the knowledge is successfully injected into the text embedding for each anatomy concept.
@@
Knowledge injection helps segmentation of long-tail classes. The 429 classes in SAT-DS-Nano follow a typical long-tail distribution, as illustrated in Fig. 8 (a) and (b). The 10 classes with maximum annotations account for 12.75% of all the annotations in SAT-DS-Nano, denoted as ‘head’ classes; On contrast, the 150 classes with minimum annotations only accounts for 3.25%, even though they comprise 34.97% of 429 classes. We compare U-Net-Ours, U-Net-CPT and U-Net-BB on the ‘head’ classes, ‘tail’ classes and the rest ones (denoted as ‘middle’ classes). In Fig. 8 (c), we observe that prompted by our proposed text-encoder, U-Net-Ours achieves the best performance on all three type of classes. On ‘head’ classes, it outperforms the second best variant by 0.71 on DSC and 2.44 on NSD. While on ‘tail’ classes, the improvement is even more significant: 1.4 on DSC and 3.23 on NSD. Meanwhile, U-Net-CPT is better than U-Net-BB on ‘tail’ but slightly underperform on ‘head’. These results suggest our proposed text encoder is consistently better in a long-tail distribution, and the knowledge injection can effectively boost the segmentation performance on ‘tail’ classes.
@@
2.3 Qualitative Results in Different Scenarios
Benefiting from the universality and text-driven features of SAT, it can be flexibly applied in clinic scenario. We verify such capability qualitatively on real clinic data. As opposed to previous universal segmentation model [46, 68, 14] that strictly relies on human interactions, SAT is built on knowledge enhanced text prompts, that can precisely understand the segmentation needs in text-form, regardless the user is a human or machine.
We utilize GPT-4 [2] to directly extract the anatomical targets of interests from real clinical report and prompt SAT to segment them on the clinical image, forming a fully automatic pipeline. In Fig. 9 and Fig. 10, we demonstrate the zero-shot performance of SAT-Pro on four cases randomly selected from clinical practice: abdominal MR examination, chest CT examination, abdominal CT examination, and lumbar spine CT examination. The basic information of patients, the report written by doctors, and targets of interests detected by GPT-4 are attached to each cases. As can be seen, the targets in reports are well detected by GPT-4 and commendably segmented by SAT-Pro: In Fig. 9, Liver, Pancreas, Kidneys, Gallbladder and Spleen are segmented in both abdominal CT and MR images; While in Fig. 10, segments of right lung, heart and bronchi are segmented in thoracic CT image, lumbar vertebrae L1 to L5 are segmented in spinal CT image. Note that in this evaluation, SAT-Pro is not fine-tuned on any data out of SAT-DS, despite the diversity of image and segmentation targets.
@@ 3 Discussion
The heterogeneity of medical images has long called for versatile, effective and flexible universal medical segmentation model to meet clinical needs. To this end, we proposed SAT, a universal medical segmentation model trained on unprecedented dataset collection, and driven by knowledge enhanced text prompt. In this section, we discuss the performance, potential and limitations of the proposed SAT, based on our experiment results in Sec. 2.
An universal model is worth 72 specialist models. Through comprehensive experiments in both region-wise and class-wise evaluations, SAT-Pro has demonstrated comparable results to an assemble of 72 nnU-Nets specialized and trained on each dataset, and even surpass them on several regions and classes. These promising results reveal universal segmentation model as a much more flexible method qualified for wide range of tasks. Even more impressive, SAT-Pro is of much smaller size, implying a more cost-effective alternative to training and deploying a series of specialist models.
Scaling models lead to better performance. Increasing model capacity has been demonstrated to effectively enhance model’s performance in large language models [110, 99], yet such scaling law remains to be verified in segmentation tasks. In this work, we build SAT-Nano (110M) and SAT-Pro (447M). In both region-wise and class-wise evaluations, SAT-Pro shows clear performance boost over SAT-Nano, outperforming the latter on most region and classes. This implies that scaling law is also applicable to universal medical segmentation, probably a promising way to continuously improve the performance.
@@
SAT can be easily transferred to clinics. Clinic utility is crucial for segmentation methods. In Sec. 2.3, we show that SAT-Pro can be applied directly to real clinical data out of the scope of SAT-DS without extra annotation and fine-tuning. Despite the diversity of image and segmentation targets from different clinic scenes, SAT can flexibly handle them with just one model. This could save laborious efforts to collect, annotate and train a bag of specialist models. More importantly, SAT can be utilized with text prompts, saving the time for per-sample human intervention, thus functioning automatically with predefined targets, reducing the cost in real use. Furthermore, as we have normalize all the image with unified voxel spacing, the successful application of SAT in clinic would pave the way for estimating the spacial size of anatomical targets directly, providing reference for analysis, diagnosis and treatment.
SAT can be a powerful agent for large language models. In Sec. 2.3, we show that SAT can be applied to segment targets extracted by GPT-4 from clinical report, providing explainable grounded report for the patients. This demonstrate the potential of SAT as a powerful agent, seamlessly integrated with any large language models, with text prompts bridging them. As grounding is an indispensable procedure in building generalist medical artificial intelligence [77], we expect the potential and significance of SAT to extend beyond a simple clinical tools.
@@
  • Firstly, the performance of SAT-Pro still lags behind nnU-Net in some region including Brain, Spine and Abdomen, and many classes especially the lesions. This implies SAT, though flexible and versatile, can not fully replace specialist models, especially in scenarios where absolute performance is critical. However, we evaluate SAT-Pro on train set, derive its averaged DSC score of all class as 83.68. We suspect its performance is still unsaturated, and continuous scaling-up universal segmentation model may close the gap and even outperforms the counterparts specialist models.
@@
  • Secondly, SAT currently only supports text as prompts, thus not intended for scenes with human interaction requirements. For example, human annotation or refinement on the segmentation results.
@@
  • Thirdly, the distribution of SAT-DS is unbalanced, leading to the limitation on capability of SAT. For instance, SAT-DS contains 189 classes for MR images, while 137 exists in ‘Brain’ region. Comprehensive annotation for MR images still lags behind CT images in the scope of public datasets.
@@
  • Fourthly, because of the imbalance between different segmentation datasets, the long-tail distribution in assembled dataset collection remain challenging for building an universal segmentation method. As demonstrated in Fig. 8 (c), the performance gap between ‘head’, ‘middle’ and ‘tail’ classes is still significant.
4 Related work
4.1 Specialist Medical Image Segmentation
Recent years have witnessed significant advancements in the filed of medical image segmentation for diverse clinical applications [84, 9, 69, 97, 5, 8, 71, 43, 70, 35, 66, 106]. Since the success of convolutional neural networks (CNNs), U-Net [92, 16, 38] have shown dominant performance for medical image segmentation, and the variants [75, 21, 57, 115, 125, 94, 119] are further developed to dive into specific clinical tasks. Meanwhile, the emerging Transformers [32, 12, 30, 112, 117, 14, 122] step in this field and obtain promising results. This stream of specialist methods typically employ a customly-designed solution and a supervised learning framework. Despite the success in specific segmentation tasks, it requires to be trained from scratch on a new segmentation task, posing limitation on the generalisability. More importantly, the data scarcity for specific tasks often significantly hamper the application in diverse and practical clinical scenarios. In this paper, we present a generalist model that covers a wide spectrum of medical image segmentation tasks.
4.2 Generalized Medical Image Segmentation
In the Medical Segmentation Decathlon [5], nnU-Net [38] takes a promising step towards generalized medical image segmentation. It offers an automatic pipeline to define the optimal network architecture and data pre-processing for each segmentation task. However, the issues of specialist model remain. It has to involve substantial overhead by training from scratch on each task, and the intrinsic correspondence among segmentation tasks is neglected. To effectively benefit a multitude of clinical segmentation tasks, advanced learning schemes have been proposed for generalized medical image segmentation, including multi-task learning [83, 120, 26, 24], transfer learning [13, 19, 104], and continual learning [118, 41, 124]. Multi-task learning incorporates multiple imaging modalities, e.g.,multi-parametric MR and multi-phase CT [83, 120], or anatomies, e.g.,multi-organ segmentation [24]. In contrast, transfer learning achieves generalized medical image segmentation by transferring segmentation model from one task to another, e.g.,from CT to MR image segmentation and vice versa [13]. Continual learning incrementally involve novel tasks into the segmentation model and avoid forgetting the old ones at the same time [41]. However, these methods require either a limited pre-defined set of tasks or substantial tuning on novel tasks, and thus can hardly scale up. The proposed method exploits textual prompt learning, i.e.,the term of the organs or lesions, to smoothly steer the segmentation model to new tasks without costly training or tuning.
4.3 Universal Medical Image Segmentation
Universal medical image segmentation paves a way to superior generality and scalability by covering a wide range of organs and lesions across various medical imaging modalities and anatomies [37]. The universal model is usually driven by visual or textual prompts to adapt to various segmentation tasks. Since the significant success of SAM [46] in computer vision, several approaches follow the interactive framework and leverage box and/or point prompts to finetune the model on medical images [68, 37, 15, 102, 27]. Other methods also exploit image examples as visual prompts [11] or integrate visual prompts with textual prompts [22]. Despite showing promising results, these semi-automated approaches rely on carefully curated visual prompts, that may face critical challenges in clinical practice. Meanwhile, the typical 2D slice-wise segmentation in most visual prompting methods [68, 37, 15] inevitably leads to suboptimal context encoding for 3D medical imaging [27]. In contrast, pure textual prompts save the effort for burdensome interaction, especially when scaling up to tens of thousands of medical images. Several works have made successful attempts to adapt textual or learnable prompts for abdominal organ and lesion segmentation [61, 114], encouraging us to take a further step towards universal medical image segmentation. However, these studies still require a pre-defined set of categories and thus suffer from the shortage of medical domain knowledge, leading to limited generality across diverse medical segmentation tasks. To bridge this gap, we explicitly leverage multimodal medical knowledge composed of anatomy atlas and definition via visual language pre-training for universal medical image segmentation.
4.4 Knowledge-enhanced Representation Learning in Medical Image Analysis
As accurate medical image analysis relies on expert domain knowledge, several efforts have attempted to involve well-established medical knowledge to strengthen deep learning methods [111, 108]. These methods can be divided into two categories, namely, model-based and learning-based methods. Specifically, the model-based methods design the network architecture based on the clinical knowledge prior [55, 28] or mimicking the clinical practice [18, 23]. The learning-based methods exploit a multi-task learning scheme that takes medical knowledge as either inputs or additional supervision [107, 123]. Additionally, in the recent literature, [52, 89, 42] train or fine-tune language models with biomedical corpus [91, 10], boosting model performance over a wide range of biomedical language tasks; while [109, 116, 54, 121] adopt knowledge-enhanced visual-language representation learning and demonstrate superior performance in downstream classification tasks. Our method aligns with the representation learning methods. While all existing methods mainly focus on incorporation with domain knowledge in text form, we explore multimodal medical knowledge alignment for both text and anatomical atlas, facilitating fine-grained visual-language knowledge injection while training universal segmentation model.
5 Dataset

### v3

@@ 2.3 Ablation Study
As will be illustrated in Sec. 4.2, to build an universal segmentation model driven by text prompts, we devise the visual encoder and decoder to process diverse medical images and perform segmentation; Meanwhile, domain knowledge is injected into the text encoder to provide precise prompts for the target of interest, i.e., the encoding of terminology. In this section, we conduct experiments and discuss the effect of different visual backbones and domain knowledge. To save computational cost, all the experiment in this section are conduct on SAT-DS-Nano dataset.
@@ 2.3.1 Effect of Visual Backbone
Extensive studies have been done discussing the applicability of different visual backbones across segmentation tasks. However, it remains unexplored on which visual backbone suits the need of universal segmentation. In addition to the ConvNet-based U-Net, we consider two alternative backbones, namely, SwinUNETR [20] and U-Mamba [18] to medical segmentation. We configure three SAT-Nano variants with SwinUNETR(107M), U-Mamba(114M) and U-Net(110M) in comparable size: For SwinUNETR, we use the same hyper-parameters as in official implementation [19]; For U-Mamba, we refer to the official implementation [53], reserving configuration of the U-Net, and only add mamba layers at the end of last 3 blocks of the U-Net encoder.
To avoid repeating multimodal knowledge injection on each visual backbones, we use MedCPT [30] for all these variants as text encoder, without loss of generality. MedCPT is a text encoder trained on 255 million in-house user click logs from PubMed [76], and shows state-of-the-art performance on various biomedical language tasks. We expect it to possess domain knowledge and embed terminologies properly as text prompts. Thus, we denote these variants as U-NET-CPT, SwinUNETR-CPT and U-Mamba-CPT. We demonstrate the region-wise evaluation results in Fig. 5. U-Net-CPT outperforms U-Mamba-CPT slightly on both DSC (0.35) and NSD (0.22) scores averaged over all classes. However, both U-Net-CPT and U-Mamba-CPT exceed SwinUNETR-CPT by a significant margin. These observations confirms our choice of U-Net as the visual backbone of SAT. For more detailed comparison on each dataset and class, we refer the reader to supplementary material Sec. 8.9 and Sec. 8.10.
@@ 2.3.2 Effect of Text Encoder
In this section, we investigate the impact of text encoder for the text-prompted universal segmentation. Specifically, we train three SAT-Nano models with three representative text encoders: the text encoder pre-trained on our multimodal medical knowledge graph, as illustrated in Sec. 4.2.2; the state-of-the-art text encoder for various medical language tasks, e.g., MedCPT; And BERT-Base, a prevalent text encoder in natural language processing but not specifically finetuned on medical corpus. For all the variants, we use U-Net as visual backbone, and denote them as U-Net-Ours, U-Net-CPT, U-Net-BB.
@@
Region-wise Evaluation. As shown in Fig. 6, the performance of U-Net-CPT and U-Net-BB are close, with U-Net-CPT slightly outperfoms U-Net-BB on both DSC (+0.43) and NSD (+0.52) scores averaged over all classes, suggesting that finetuned on medical corpus, MedCPT might hold minor advantage. By contrast, U-Net-Ours surpasses U-Net-CPT consistently on all regions and lesions, leading notable margins on both DSC (+1.54) and NSD (+2.65) scores after average over all classes. This demonstrates the effectiveness of our proposed multimodal knowledge injection. For more detailed comparison on each dataset and class, we refer the reader to supplementary material Sec. 8.7 and Sec. 8.8.
@@
Medical Concept-to-definition Retrieval. To demonstrates advantage our proposed knowledge injection on capturing the critical features of medical terminologies, we also evaluate the three text encoders on concept to definition retrieval task, with the human anatomy knowledge we collect, totaling 6,502 anatomy concepts and definition pairs. We find that the Recall1 (R1) for BERT-Base is merely 0.08%, suggesting it can hardly understand these anatomy concepts or possess almost no domain knowledge; The R1 for MedCPT is 11.19%. Though this is a significant improvement over BERT-Base, MedCPT still struggles to distinguish these concepts; By contrast, our proposed text encoder get 99.18% R1, indicating the knowledge is successfully injected into the text embedding for each anatomy concept.
@@
Knowledge Injection Helps Segmentation of Long-tail Classes. The 429 classes in SAT-DS-Nano follow a typical long-tail distribution, as illustrated in Fig. 7 (a) and (b). The 10 classes with maximum annotations account for 12.75% of all the annotations in SAT-DS-Nano, denoted as ‘head’ classes; On contrast, the 150 classes with minimum annotations only accounts for 3.25%, even though they comprise 34.97% of 429 classes. We compare U-Net-Ours, U-Net-CPT and U-Net-BB on the ‘head’ classes, ‘tail’ classes and the rest ones (denoted as ‘middle’ classes). In Fig. 7 (c), we observe that prompted by our proposed text-encoder, U-Net-Ours achieves the best performance on all three type of classes. On ‘head’ classes, it outperforms the second best variant by 0.71 on DSC and 2.44 on NSD. While on ‘tail’ classes, the improvement is even more significant: 1.4 on DSC and 3.23 on NSD. Meanwhile, U-Net-CPT is better than U-Net-BB on ‘tail’ but slightly underperform on ‘head’. These results suggest that our proposed knowledge injection can effectively boost the segmentation performance on ‘tail’ classes.
@@
2.4 Qualitative Results – SAT as an Interface Between Language and Segmentation
Benefiting from the universality and text-driven features of SAT, it can be flexibly applied as an interface between natural language and segmentation. Specifically, we utilize GPT-4 [2] to directly extract the anatomical targets of interests from real clinical report and prompt SAT to segment them on the clinical image. In Fig. 8 and Fig. 9, we demonstrate the zero-shot performance of SAT-Pro on four cases randomly selected from clinical practice: abdominal MR examination, chest CT examination, abdominal CT examination, and lumbar spine CT examination. The basic information of patients, the report written by doctors, and targets of interests detected by GPT-4 are attached to each cases. As can be seen, the targets in reports can be well detected by language model (GPT-4) and commendably segmented by SAT-Pro: In Fig. 8, Liver, Pancreas, Kidneys, Gallbladder and Spleen are segmented in both abdominal CT and MR images; While in Fig. 9, segments of right lung, heart and bronchi are segmented in thoracic CT image, lumbar vertebrae L1 to L5 are segmented in spinal CT image.
@@ 3 Discussion
Developing specialist models on individual tasks has been the dominant solution for medical image segmentation for years [27, 15, 44, 95, 102, 79, 99, 21, 11, 19, 93, 98, 12, 101]. Yet the heterogeneity of medical images has long called for effective and flexible universal medical segmentation model to meet clinical needs. To this end, we proposed a universal segmentation model, trained on unprecedented dataset collection, and driven by knowledge enhanced text prompt. In this section, we discuss the significance, potential and limitations of the proposed SAT, based on our experiment results in Sec. 2.
A Universal Model is Worth 72 Specialist Models. Through comprehensive internal evaluation from several aspects, SAT-Pro has demonstrated comparable results to an assemble of 72 specialized nnU-Nets, and even surpass them on lots of regions and classes. While on external evaluation, SAT-Nano exhibits superior generalization ability than nnU-Nets. Therefore, despite the diversity of image and segmentation targets from different clinic scenes, SAT can flexibly handle them with just one model. SAT along with the unprecedented dataset collection and standardisation procedure, represent an important step forward for universal segmentation model in medical scenario.
SAT as a Strong Pre-trained Segmentation Model. When fine-tuned on specific dataset, SAT-Pro-Ft can further improve the performance, achieving superior results over the counterpart nnU-Nets. Thus, it naturally acts a foundational segmentation model, pre-trained on large-scale and diverse data, holding potentials to be adapted to specific tasks, balancing the clinical requirements between universal solution and specialist model.
Text As Prompt for Segmentation. As an automatic method prompted by text, SAT outlines a new path apart from recent works in interactive universal segmentation [54, 17, 87, 87]. In both internal and external evaluation, SAT saves time from laborious spatial prompting, showing superior scalability in deployment, provides semantic and robust results regardless of manual errors in drawing box prompts.
SAT Can Be Used As An Agent to Bridge Language and Segmentation. In Sec. 2.4, we show that SAT can be applied to segment targets extracted by language model from clinical report, providing explainable grounded report in clinical scenarios. This demonstrate the potential of SAT as a powerful out-of-the-box agent, seamlessly integrated into any large language models via text prompts. As grounding is an indispensable procedure for building generalist medical artificial intelligence [63], we expect the potential and significance of SAT to extend beyond a simple clinical tools, for example, providing grounding annotations for medical datasets in a scalable way [97].
Scaling Models Lead to Better Performance. Increasing model capacity has been demonstrated to effectively enhance model’s performance in large language models [92, 84], yet such scaling law remains to be verified in segmentation tasks. In this work, we build SAT-Nano (110M) and SAT-Pro (447M). In both region-wise and class-wise evaluations, SAT-Pro shows clear performance boost over SAT-Nano, outperforming the latter on most region and classes. This implies that scaling law is also applicable to universal medical segmentation, indicating a promising way to continuously improve the performance.
@@
  • The performance of SAT-Pro still lags behind nnU-Nets in some region including Brain, Spine and Abdomen, and many classes. In future work, we will continue scaling up model size, to further close the gap and even outperforms the counterparts specialist models.
@@
  • SAT currently only supports text as prompts, thus not intended for scenes with human interaction requirements. For example, manual refinement on the segmentation results. We plan to incorporate more interaction interfaces, for example, boxes, scribbles.
@@
  • The distribution of SAT-DS is unbalanced, leading to the limitation on capability of SAT. For instance, SAT-DS contains 189 classes for MR images, while 137 classes are on ‘Brain’ region. Comprehensive annotation for MR images still lags behind CT images in the scope of public datasets.
@@
  • Due to the imbalance between different segmentation datasets, the long-tail distribution in assembled dataset collection remain challenging for building an universal segmentation method. As demonstrated in Fig. 7 (c), the performance gap between ‘head’, ‘middle’ and ‘tail’ classes is still significant.
4 Method
4.1 Dataset Construction